### v1

@@ 3.2 Turing Machines and the Halting Problem
Throughout this review, we will see that most of the undecidability results in physics rely on an embedding of a Turing Machine (see Fig. 1) into a physical system. Let us now introduce the formal definition of a Turing machine. There are several equivalent options for that. We will take the definition from E. Bernstein (1997). For other options, we refer for instance to Sipser (2012).
@@
It is not difficult to see that, given a TM $T$ and an input $x$, we can define a new Turing machine $T^{\prime}$ which halts on a blank input if and only it $T$ halts on $x$ and, in that case, both give the same output. $T^{\prime}$ is just the machine that, on a blank input, first writes $x$ on the tape and then runs as machine $T$.
It is a remarkable fact, already proven by Turing (1936) that there exist universal Turing machines. A universal TM is a Turing machine that can simulate the behavior of any other TM in the sense that, if given as input the description of a Turing machine $T$, it halts if and only if $T$ halts on the blank input and, in that case, its output is identical to that of $T$. This implies that a universal TM can simulate the behavior of any other TM initialized in any possible input.
@@
A function $f:\tilde{\Sigma}^{*}\rightarrow\Sigma^{*}$ is computable if there exists a TM with alphabet $\Sigma$ which, on input $x\in\tilde{\Sigma}^{*}$, halts and leaves as output $f(x)$. A function is uncomputable if it is not computable.
@@ 3.2.1 The Halting Problem
    If $H(x)$ outputs YES, then $D(x)$ enters an infinite loop.
@@
    If $H(x)$ outputs NO, then $D(x)$ halts.
@@
Therefore, our assumption that $H$ exists must be false, and the halting problem is undecidable. For a formal proof see Hopcroft et al. (2006); Sipser (2012).
@@ 3.2.2 Church’s Theorem
In section 2.2 we have mentioned that, in addition to Turing, also A. Church found a negative answer to the decision problem with his $\lambda$-calculus. In his 1936 paper “An Unsolvable Problem of Elementary Number Theory” he first provided examples in elementary number theory and topology, and later proved several undecidability theorems in problems with integer recursive functions. (All these results are usually referred together as “Church’s Theorem” Church (1936)).
Although Church’s result preceded Turing’s, it is not usually mentioned as an undecidability result together with Gödel’s and Turing’s, probably because Church and Turing quickly demonstrated jointly that $\lambda$-calculus and Turing machines were equivalent, and the latter was developed as a description of a real computational machine, while the mathematical description of the former was more abstract.
However, we mention it here because Church’s results were used by Shapiro (1956) to prove undecidability of other important decision problems in integer recursive functions, and in turn, one of those results was used by Komar (1964) to obtain an undecidable problem in quantum field theory (see section 4).
@@ 3.3 Relation between (Turing) undecidability and (Gödel) axiomatic independence
In some sense, (Turing) undecidability implies (Gödel) axiomatic independence. In this brief section, we will sketch this connection following Poonen (2014).
Consider a formal system, which is effectively axiomatized, and a decision problem with (infinitely many) instances $i$. As it is usually the case, assume that for each instance of the decision problem, one can construct (algorithmically) a statement $Y_{i}$ so that if $Y_{i}$ is provable within the formal system, then the answer to instance $i$ is YES, and if its negation $\neg Y_{i}$ is provable, then the answer to instance $i$ is NO.
@@
As mentioned above, we will mainly focus on (Turing) undecidability. But using this connection, one obtains (Gödel) axiomatic independence for each one of the undecidable problems reviewed here. Moreover, if the undecidability originates from the halting problem, for any given consistent formal system one can use this connection in order to pinpoint an individual instance of the decision problem that cannot be decided on the basis of the axioms. The reason is that, by Gödel’s second incompleteness theorem, consistency of a formal system cannot be proven within that system (unless it is inconsistent). Hence, we could, at least in principle, set up an explicit Turing machine that recursively lists all theorems that are provable from the axioms and halts iff a given inconsistent statement like “$0=1$” appears. If the formal system is consistent, then this Turing machine will never halt, but this fact will not be provable within the formal system. For ZFC, the axiomatic foundation of set theory and the most commonly used foundation of mathematics, this has been worked out by Aaronson and Yedidia (2016) and improved by O’Rear (2016).
@@ 3.4.1 Diophantine Equations and Hilbert’s Tenth Problem
Making an actual theorem out of this conjecture turned out to be an arduous task, and it did not occur until 1995 when Andrew Wiles presented his celebrated proof Wiles (1995); Taylor and Wiles (1995). In general, the great difficulty of proving the existence of solutions for the various types of Diophantine equations led Hilbert to include this problem in his famous 1900 list of 23 mathematical problems for the 20th century.
@@
In 1971 Matiyasevich (1993), using previous results by Davis, Putnam and Robinson (Davis et al. (1961)), proved that such an algorithm does not exist, i.e. the existence of integer solutions for the Diophantine equations is undecidable.
This important result in mathematics has been used to prove undecidable questions in different physical systems, such as integrability of classical Hamiltonians or the existence of chaos in dynamical systems (da Costa and Doria (1991a, b)); and, more recently, solvability of quantum control problems (Bondar and Pechen (2020)), and supersymmetry breaking in quantum field theory models (Tachikawa (2023)) (sections 4, 5.2.1, and 5.4, respectively).
@@ 3.4.2 Post Correspondence Problem
The Post Correspondence Problem (PCP) was introduced by Post (1946) and is very often used in proofs of undecidability because it has a simpler structure than the halting problem. We follow Sipser (2012), but this is also covered in Hopcroft et al. (2006), and (Goodman-Strauss (2010)) for a more informal and didactic introduction.
@@
By a reduction to the Turing Machine halting problem, E. Post proved that the PCP is undecidable. The PCP was used as the basis to prove the undecidability of the Matrix Mortality problem that we comment on next.
@@ 3.4.3 Matrix mortality
The Matrix Mortality Problem was proven to be undecidable even with $d=3$ and $n=16$ by Paterson (1970) with a reduction to the PCP. Improved pairs $(d,n)$ for which the Matrix mortality problem remains undecidable were obtained later by several authors (see for instance the review Blondel and Tsitsiklis (2000)).
This problem has been used to prove undecidability results in tensor networks (Kliesch et al. (2014); De las Cuevas et al. (2016)) and quantum measurement theory (Eisert et al. (2012)) that are discussed in sections 5.1.3 and 5.2.1 respectively.
@@ 3.4.4 Word Problem
In the 50s, Novikov (1955) and Boone (1959) independently proved that there exist finitely presented groups for which the Word Problem is undecidable.
Several of the works we will review here, like Geroch and Hartle (1986); Slofstra (2020, 2019); Fritz (2021); Mančinska and Roberson (2020); Fu et al. (2021), are based on the undecidability of the Word Problem.
@@ 3.4.5 Wang tiles
In 1961 H. Wang conjectured that a finite set of tiles can tile the entire plane if and only if it can tile the plane periodically (Wang (1961)). A corollary of this is that the decision problem of determining whether or not a given set of Wang tiles can tile the plane would be decidable. However, in 1966 his student R. Berger disproved the conjecture by embedding a Turing Machine into a tile set in such a way that they can tile the plane if and only if the machine will never halt. Therefore the Wang tiling problem is undecidable. Berger’s construction was substantially simplified by Robinson (1971).
With this new universal system at their disposal, researchers tried to spread undecidability by making different kinds of connections, with cellular automata being the first candidates due to their similar structure when mapping their temporal dimension to one of the spatial dimensions in the tilings (see next section). The regular structure of tilings in the plane allows as well for direct mappings to different physical systems, to prove undecidable properties in them. The results by Kanter (1990) (section 4), Cubitt et al. (2015b) and Scarpa et al. (2020) (section 5), are based on or make use of mappings between tilings and 2-dimensional Hamiltonians.
The classical reference for tilings is Grünbaum and Shephard (1986). See as well Goodman-Strauss (2010) for an informal and didactic introduction to undecidability in tiling problems.
@@ 3.4.6 Cellular Automata
Conway’s Game of Life was developed by mathematician John Conway in 1970 and became known thanks to a Scientific American article (Gardner (1970)). From a finite initial condition in the center of a 2D grid, this CA with two states (alive and dead) and first neighbor interactions (range $r=1$) can evolve into very diverse complex behavior, including universal computation (Rendell (2015); Blundell (2017)).
But there are even simpler examples with this universal capability, such as rule 110: the simplest so far (1D, two-states, $r=1$). This CA was conjectured to be Turing complete in 1985 by S. Wolfram and proved to be so in 2004 by emulating cyclic tag systems, in turn proven universal by embedding a Turing Machine into them (Cook (2004)).
Stephen Wolfram has been a prominent contributor to the development and study of CAs since the 1980s (Wolfram (1984, 1985)). His book “A New Kind of Science” Wolfram (2002) discusses various results about universal computation and undecidability in these and other systems, such as universality of rule 110, finding a universal Turing machine with $2$ internal states and an alphabet of size $5$, and proposing an even simpler candidate with $2$ internal states and $3$ alphabet symbols (that eventually was proved universal by an undergraduate student).
J. Kari has studied undecidable properties of CA using mappings to Wang tiles to prove general results such as Rice’s theorem for CAs (Kari (1994)). The application is restricted to limit sets, which are configurations that can occur after arbitrary many computational steps and therefore represent long-term behaviour. Kari (2012) reviews the main results about undecidability in CAs considering diverse properties including reversibility, time-symmetry, conservation laws, chaotic dynamics, and the influence of the dimension in undecidability.
It is precisely this diversity, together with their local interactions and their regularity in space, that make CAs well-suited for embedding into planar physical systems, in which the temporal evolution of a 1D CA is mapped to a spatial dimension of the ground state of a 2D Ising lattice, for instance. We will discuss results about undecidability in physical systems obtained in this way in the next sections (Omohundro (1984); Gu et al. (2009); Gu and Perales (2012)).
Let us finally point out that Israeli and Goldenfeld (2004, 2006) constructed local coarse-grained descriptions of 1D CAs that can emulate the large-scale behavior of the original systems without accounting for small-scale details. They showed that although some of the original CAs present undecidable properties, the coarse-grained large-scale construction can be predictable and even decidable at some level of description.
@@
There are undecidable questions in mathematics, the language of physics is mathematics, therefore it is possible that there are undecidable questions in physics. The connections of this simple argument with the strange uncertainties provided by quantum mechanics at the beginning of the 20th century were soon noticed by some, but it seems that Gödel himself did not want to explore this path [^15] John Wheeler asked Gödel about the connection between his incompleteness and Heisenberg’s Uncertainty Principle and “Gödel got angry and threw me out of his office” Barrow (2011)..
We can consider the very Turing Machine and its halting problem as the first result of undecidability in the physical world. Except for the idealization of the infinite tape (see section 6.2), it is a mechanical machine that can be perfectly built. However, Alan Turing’s motivation was not to consider that connection either, but to rigorously establish the concept of algorithm and universal machine (and as a consequence lay the foundations of theoretical computer science).
@@
Among the former, the philosopher of science Karl Popper (Popper (1950)) used Gödel’s results to describe a mechanical machine called ‘predictor’ with a ‘calculator’ [^16] With remarkable similarities to a Turing Machine: “A classical mechanical calculating and predicting machine which is so constructed as to produce permanent records of some kind (such as a tape with holes punched into it) capable of being interpreted as predictions of the positions, velocities, and masses of physical particles. It will be shown that such a machine can never fully predict every one of its own future states …Moreover, can never fully predict, or be predicted by, any sufficiently similar machine with which it interacts.” (Popper (1950)) to hypothesize that “most systems of physics, including classical physics and quantum physics, are indeterministic in perhaps an even more fundamental sense than the one usually ascribed to the indeterminism of quantum physics”. Schlegel (1967) treated incompleteness in all sciences in the light of Gödel’s theorems (in particular uncertainty in quantum mechanics) and Zwick (1978) and Rothstein (1982) explored the connections with the measurement problem and irreversibility, respectively. None of them included mathematical proofs.
@@
The first rigorous result in the field was obtained by Komar (1964), proving that the macroscopic distinguishability of two arbitrary states with infinite degrees of freedom in quantum field theory is undecidable. He mapped this system into a decision problem about distinguishability of two infinite sequences of integers that was proven undecidable by Shapiro (1956) via a reduction to Church’s theorem (section 3.2.2).
@@
The next batch of results were obtained two decades later, in the 1980s and 1990s. Pour-El and Richards (1981) found non-computable weak solutions for a particular three-dimensional wave equation with computable initial conditions. Fredkin and Toffoli (1982) managed to construct a reversible “billiard ball” model of universal computation with hard frictionless balls that could perform all necessary logic operations by elastic collisions [^17] The authors focus on the computational capabilities of their model, so do not mention the undecidable questions inherent to any universal system of computation, as first shown by Turing. Our interest lies in the fact that they prove that a simple physical system, formally identical to the atomic model of the classical kinetic theory of perfect gases, can perform universal computation and therefore present undecidable properties.. Omohundro (1984) simulated an arbitrary CA with a system of partial differential equations, proving therefore that this system is able to perform universal computation as well (see footnote 17).
@@
In a work aiming to find a criterion to establish whether a physical theory should be acceptable by distinguishing between computable and measurable quantities, Geroch and Hartle (1986) presented as an example an observable in quantum gravity for closed cosmologies –formulated as a sum over histories–, whose expectation value is not computable [^18] The authors recognized that the mathematics of quantum gravity had not yet been developed to the point to have a complete mathematical formulation of what the equation providing the expectation value meant.. They proposed that the calculation should involve the question of whether two simplicial 4-manifolds are topologically identical, which was proved undecidable by Haken (1973) via a reduction to the word problem (section 3.4.4). It is interesting to note that this example introduces infinite magnitudes in a natural way since the calculation of the expected value involves the infinite sum over histories (see section 6.2).
@@
We have seen in section 3.4.5) that tiling problems are very natural candidates to transfer undecidability to physical systems. This was the approach that Kanter (1990) took to prove undecidable properties in a 1D spin system (anisotropic Potts Hamiltonian) using a map to the ’domino snake’ problem, which is a connectivity variation of a Wang tiling problem. The author also indicated how his arguments could be extended to higher dimensions, dynamical evolution, other spin systems, and to some spatial and temporal correlation functions in these systems. He also introduced an uncertainty principle for classical systems caused by undecidability [^19] To our knowledge no further work was done in these directions..
@@
Around the same time, Moore (1990, 1991) showed how to embed a universal Turing Machine into a continuous dynamical system of just three degrees of freedom, corresponding to the motion of a single particle in a finite-sized, three-dimensional box. His construction uses plane and parabolic mirrors in the box to modify the chaotic trajectories of the dynamical system in order to obtain the Turing Machine operations needed for universality. Virtually any question about its long-term behavior is undecidable. We refer to Bennett (1990, 1995) for a discussion of this remarkable achievement.
@@
Taking a very different approach and mapping, da Costa and Doria (1991a, b) (see as well Chaitin et al. (2011)) encode Diophantine equations in classical mechanics and prove undecidability of questions about integrability of certain Hamiltonians and about the existence of chaos in dynamical systems. They rely on results about undecidability of the existence of roots in certain classes of real functions, obtained in turn by mapping to Diophantine equations (Richardson (1968); Wang (1974)). But as the authors themselves write, “While we have given general existential proofs concerning undecidability and incompleteness in physics, we have not considered up to now the same questions with respect to particular physically interesting systems …dealing with specific situations is assuredly much more difficult than to prove general facts”. One can find an introduction to these results in Stewart (1991). Calude et al. (1995) use the same tools to provide a couple of undecidability results about integral solutions of the one-dimensional heat equation and the electrostatic plane problem.
@@
da Costa et al. (1990) obtained undecidability results in the sense of Gödel in general relativity. They constructed an axiomatic framework for it within Zermelo–Fraenkel set theory and showed that some statements are axiomatically independent. See also Doria and Doria (2011); Chaitin et al. (2011).
@@
Finally, Lloyd (1993, 1994) pioneered fruitful interconnections between quantum information theory and undecidability, which blossomed more than 15 years later, as we will review in detail in the next section. In Lloyd (1993, 1994), it is argued that certain spectral properties of quantum mechanical operators are undecidable. The argument is based on the following idea: the unitary evolution $U$ associated to the time evolution of a universal quantum computer has invariant subspaces with discrete spectrum, corresponding to computations that halt, and other invariant subspaces with continuous spectrum, corresponding to computations that do not halt. Via the undecidability of the halting problem, Lloyd concludes that it is undecidable to know if the quantum state associated to a given program (in the infinite-dimensional space in which $U$ is defined) has overlap, or not, with an invariant subspace having discrete spectrum [^20] See Lloyd (2016); Cubitt et al. (2016) for a discussion on the relation between Lloyd (1993, 1994) and Cubitt et al. (2015b, 2022)..

### v2

@@ 3.2 Turing Machines and the Halting Problem
Throughout this review, we will see that most of the undecidability results in physics rely on an embedding of a Turing Machine into a physical system (see Fig. 1).. Let us now introduce the formal definition of a Turing machine. There are several equivalent options for that. We will take the definition from [52]. For other options, we refer for instance to [53].
@@
It is not difficult to see that, given a TM $T$ and an input $x$, we can define a new Turing machine $T^{\prime}$ which halts on a blank input if and only if $T$ halts on $x$ and, in that case, both give the same output. $T^{\prime}$ is just the machine that, on a blank input, first writes $x$ on the tape and then runs as machine $T$.
It is a remarkable fact, already proven by Turing in [30] that there exist universal Turing machines. A universal TM is a Turing machine that can simulate the behavior of any other TM in the sense that, if given as input the description of a Turing machine $T$, it halts if and only if $T$ halts on the blank input and, in that case, its output is identical to that of $T$. This implies that a universal TM can simulate the behavior of any other TM initialized in any possible input.
@@
A function $f:\tilde{\Sigma}^{*}\to\tilde{\Sigma}^{*}$ is computable if there exists a TM with alphabet $\Sigma$ which, on input $x\in\tilde{\Sigma}^{*}$, halts and leaves as output $f(x)$. A function is uncomputable if it is not computable.
@@
Note that, given a decision function, the set of all elements in $\tilde{\Sigma}^{*}$ whose output is YES gives a one-to-one correspondence between decision functions and subsets of $\tilde{\Sigma}^{*}$, usually called languages. Using this correspondence, the problem of determining the output of a decision function is the same as the problem of determining whether a given string in $\tilde{\Sigma}^{*}$ belongs to a language.
@@ 3.2.1 The Halting Problem
    If $H(x,x)$ outputs YES, then $D(x)$ enters an infinite loop.
@@
    If $H(x,x)$ outputs NO, then $D(x)$ halts.
@@
Therefore, our assumption that $H$ exists must be false, and the halting problem is undecidable. For a formal proof see [36, 53].
@@ 3.2.2 Church’s Theorem
In section 2.2 we have mentioned that, in addition to Turing, also Church found a negative answer to the decision problem with his $\lambda$-calculus. In his 1936 paper “An Unsolvable Problem of Elementary Number Theory” he first provided examples in elementary number theory and topology, and later proved several undecidability theorems in problems with integer recursive functions. (All these results are usually referred together as “Church’s Theorem” [29]).
Although Church’s result preceded Turing’s, it is less frequently cited than Gödel’s and Turing’s, probably because Church and Turing quickly demonstrated jointly that $\lambda$-calculus and Turing machines were equivalent, and the latter was developed as a description of a real computational machine, while the mathematical description of the former was more abstract.
However, we mention it here because Church’s results were used by Shapiro [54] to prove undecidability of other important decision problems in integer recursive functions, and in turn, one of those results was used by Komar [3] to obtain an undecidable problem in quantum field theory (see section 4).
@@ 3.3 Relation between (Turing) undecidability and (Gödel) axiomatic independence
In some sense, (Turing) undecidability implies (Gödel) axiomatic independence. In this brief section, we will sketch this connection following the review by Poonen [23].
Consider a formal system, which is effectively axiomatized, and a decision problem with (infinitely many) instances $i$. As is usually the case, assume that for each instance of the decision problem, one can construct (algorithmically) a statement $Y_{i}$ so that if $Y_{i}$ is provable within the formal system, then the answer to instance $i$ is YES, and if its negation $\neg Y_{i}$ is provable, then the answer to instance $i$ is NO.
@@
As mentioned above, we will mainly focus on (Turing) undecidability. But using this connection, one obtains (Gödel) axiomatic independence for each one of the undecidable problems reviewed here. Moreover, if the undecidability originates from the halting problem, for any given consistent formal system one can use this connection in order to pinpoint an individual instance of the decision problem that cannot be decided on the basis of the axioms. The reason is that, by Gödel’s second incompleteness theorem, consistency of a formal system cannot be proven within that system (unless it is inconsistent). Hence, we could, at least in principle, set up an explicit Turing machine that recursively lists all theorems that are provable from the axioms and halts iff a given inconsistent statement like “$0=1$” appears. If the formal system is consistent, [^13] Technically, a somewhat stronger assumption than consistency is needed here to make this argument go through; namely, that the system cannot prove any false arithmetic statements. “Rosser’s trick” and related methods allow these types of argument to be strengthened so that they do not require any such additional assumptions, see e.g. [55]. then this Turing machine will never halt, but this fact will not be provable within the formal system. For Zermelo-Fraenkel set theory with the Axiom of Choice (ZFC), the axiomatic foundation of set theory and the most commonly used foundation of mathematics, this has been worked out by Aaronson and Yedidia (2016) [56] with an explicit description of a 7910-state Turing machine with one tape and a two-symbol alphabet that cannot be proved to run forever in ZFC, and improved by O’Rear (2016) [57].
@@ 3.4.1 Diophantine Equations and Hilbert’s Tenth Problem
Making an actual theorem out of this conjecture turned out to be an arduous task, and it did not occur until 1995 when Andrew Wiles presented his celebrated proof [58, 59]. In general, the great difficulty of proving the existence of solutions for the various types of Diophantine equations led Hilbert to include this problem in his famous 1900 list of 23 mathematical problems for the 20th century.
@@
In 1971 Matiyasevich [60], using previous results by Davis, Putnam and Robinson [61], proved that such an algorithm does not exist, i.e. the existence of integer solutions for the Diophantine equations is undecidable.
This important result in mathematics has been used to prove undecidable questions in different physical systems, such as integrability of classical Hamiltonians or the existence of chaos in dynamical systems (da Costa and Doria (1991) [62, 25]); and, more recently, solvability of quantum control problems (Bondar and Pechen (2020) [63]), and supersymmetry breaking in quantum field theory models (Tachikawa (2023) [64]) (sections 4, 5.2.1, and 5.4, respectively).
@@ 3.4.2 Post Correspondence Problem
The Post Correspondence Problem (PCP) was introduced by E. Post in 1946 [65] and is very often used in proofs of undecidability because it has a simpler structure than the halting problem. We follow [53], but this is also covered in [36], and [24] for a more informal and didactic introduction.
@@
By a reduction to the Turing Machine halting problem, E. Post proved that the PCP is undecidable. The PCP was used as the basis to prove the undecidability of the Matrix Mortality problem which we comment below.
@@ 3.4.3 Matrix mortality
The Matrix Mortality Problem was proven to be undecidable even with $d=3$ and $n=16$ by Paterson in 1970 [66] with a reduction to the PCP. Improved pairs $(d,n)$ for which the Matrix mortality problem remains undecidable were obtained later by several authors [67].
This problem has been used to prove undecidability results in tensor networks (Kliesch et al. (2014) [68], De las Cuevas et al. (2016) [69]) and quantum measurement theory (Eisert (2012) [70]) that are discussed in sections 5.1.3 and 5.2.1 respectively.
@@ 3.4.4 Word Problem
In the 50s, Novikov [71] and Boone [72] independently proved that there exist finitely presented groups for which the Word Problem is undecidable.
Several of the works we will review here, like the ones of Geroch and Hartle (1986) [4], Morton and Biamonte (2012) [73], Slofstra (2019, 2020) [74, 75], Fritz (2021) [76], Mancinska and Roberson (2020) [77] and Fu et al.(2021) [78], are based on the undecidability of the Word Problem.
@@ 3.4.5 Wang tiles
In 1961 H. Wang conjectured that a finite set of tiles can tile the entire plane if and only if it can tile the plane periodically ([79]). A corollary of this is that the decision problem of determining whether or not a given set of Wang tiles can tile the plane would be decidable. However, in 1966 his student R. Berger disproved the conjecture by embedding a Turing Machine into a tile set in such a way that they can tile the plane if and only if the machine will never halt. Therefore the Wang tiling problem is undecidable. Berger’s construction was substantially simplified by Robinson (1971) [80].
With this new universal system at their disposal, researchers tried to spread undecidability by making different kinds of connections, with cellular automata being the first candidates due to their similar structure when mapping their temporal dimension to one of the spatial dimensions in the tilings (see next section). The regular structure of tilings in the plane allows as well for direct mappings to different physical systems, to prove undecidable properties in them. The results by Kanter (1990) [81] (section 4), Cubitt et al. (2015) [82] and Scarpa et al. (2020) [83] (section 5), are based on or make use of mappings between tilings and 2-dimensional Hamiltonians.
The classical reference for tilings is the book by Grünbaum and Shephard [84]. See as well [24] for an informal and didactic introduction to undecidability in tiling problems.
@@ 3.4.6 Cellular Automata
Conway’s Game of Life was developed by mathematician John Conway in 1970 and became known thanks to a Scientific American article [85]. From a finite initial condition in the center of a 2D grid, this CA with two states (alive and dead) and first neighbor interactions (range $r=1$) can evolve into very diverse complex behavior, including universal computation [86, 87].
But there are even simpler examples with this universal capability, such as rule 110: the simplest so far (1D, two-states, $r=1$). This CA was conjectured to be Turing complete in 1985 by S. Wolfram and proved to be so in 2004 by emulating cyclic tag systems, in turn proven universal by embedding a Turing Machine into them [88].
Stephen Wolfram has been a prominent contributor to the development and study of CAs since the 1980s [89, 90]. His book “A New Kind of Science” [91] discusses various results about universal computation and undecidability in these and other systems, such as universality of rule 110, finding a universal Turing machine with $2$ internal states and an alphabet of size $5$, and proposing an even simpler candidate with $2$ internal states and $3$ alphabet symbols (that eventually was proved universal by the undergraduate student A. Smith in 2007 [92]).
J. Kari [93] has studied undecidable properties of CA using mappings to Wang tiles to prove general results such as Rice’s theorem for CAs. The application is restricted to limit sets, which are configurations that can occur after arbitrary many computational steps and therefore represent long-term behaviour. In [94] Kari reviews the main results about undecidability in CAs considering diverse properties including reversibility, time-symmetry, conservation laws, chaotic dynamics, and the influence of the dimension in undecidability.
It is precisely this diversity, together with their local interactions and their regularity in space, that make CAs well-suited for embedding into planar physical systems, in which the temporal evolution of a 1D CA is mapped to a spatial dimension of the ground state of a 2D Ising lattice, for instance. We will discuss results about undecidability in physical systems obtained in this way in the next sections (Omohundro (1984) [95], Gu et al. (2009) [96], Gu and Perales (2012) [97]).
Let us finally point out that Israeli and Goldenfeld [98, 99] constructed local coarse-grained descriptions of 1D CAs that can emulate the large-scale behavior of the original systems without accounting for small-scale details. They showed that although some of the original CAs present undecidable properties, the coarse-grained large-scale construction can be predictable and even decidable at some level of description.
@@
There are undecidable questions in mathematics, the language of physics is mathematics, therefore it is possible that there are undecidable questions in physics. The conceptual similarities of this simple argument with the strange uncertainties provided by quantum mechanics at the beginning of the 20th century were soon noticed by some, but it seems that Gödel himself did not want to explore this path [^16] John Wheeler asked Gödel about the connection between his incompleteness and Heisenberg’s Uncertainty Principle and “Gödel got angry and threw me out of his office” [15]..
We can consider the very Turing Machine and its halting problem as the first result of undecidability in the physical world. Except for the idealization of the infinite tape (see section 6.2), it is a mechanical machine that can be perfectly built. However, Alan Turing’s motivation was not to consider that connection either, but to rigorously establish the concept of an algorithm and universal machine (and as a consequence lay the foundations of theoretical computer science).
@@
Among the former, the philosopher of science Karl Popper [100] used Gödel’s results to describe a mechanical machine called ‘predictor’ with a ‘calculator’ [^17] With remarkable similarities to a Turing Machine: “A classical mechanical calculating and predicting machine which is so constructed as to produce permanent records of some kind (such as a tape with holes punched into it) capable of being interpreted as predictions of the positions, velocities, and masses of physical particles. It will be shown that such a machine can never fully predict every one of its own future states …Moreover, can never fully predict, or be predicted by, any sufficiently similar machine with which it interacts.” [100] to hypothesize that “most systems of physics, including classical physics and quantum physics, are indeterministic in perhaps an even more fundamental sense than the one usually ascribed to the indeterminism of quantum physics”. Schlegel (1967) [101] treated incompleteness in all sciences in the light of Gödel’s theorems (in particular uncertainty in quantum mechanics) and Zwick (1978) [102] and Rothstein (1982) [103] explored the connections with the measurement problem and irreversibility, respectively. None of them included mathematical proofs.
@@
The first rigorous result in the field was obtained by Komar (1964) [3], proving that the macroscopic distinguishability of two arbitrary states with infinite degrees of freedom in quantum field theory is undecidable. He mapped this system into a decision problem about distinguishability of two infinite sequences of integers that was proven undecidable via a reduction to Church’s theorem [54] (section 3.2.2).
@@
The next batch of results were obtained two decades later, in the 1980s and 1990s. Pour-El and Richards (1981) [104] found non-computable weak solutions for a particular three-dimensional wave equation with computable initial conditions. Fredkin and Toffoli (1982) [105] managed to construct a reversible “billiard ball” model of universal computation with hard frictionless balls that could perform all necessary logic operations by elastic collisions [^18] The authors focus on the computational capabilities of their model, so do not mention the undecidable questions inherent to any universal system of computation, as first shown by Turing. Our interest lies in the fact that they prove that a simple physical system, formally identical to the atomic model of the classical kinetic theory of perfect gases, can perform universal computation and therefore present undecidable properties.. Omohundro (1984) [95] simulated an arbitrary CA with a system of partial differential equations, proving therefore that this system is able to perform universal computation as well (see footnote 18).
@@
In a work aiming to find a criterion to establish whether a physical theory should be acceptable by distinguishing between computable and measurable quantities, Geroch and Hartle (1986) [4] presented as an example an observable in quantum gravity for closed cosmologies –formulated as a sum over histories–, whose expectation value is not computable [^19] The authors recognized that the mathematics of quantum gravity had not yet been developed to the point to have a complete mathematical formulation of what the equation providing the expectation value meant.. They proposed that the calculation should involve the question of whether two simplicial 4-manifolds are topologically identical, which was proved undecidable via a reduction to the word problem [106] (section 3.4.4). It is interesting to note that this example introduces infinite magnitudes in a natural way since the calculation of the expected value involves the infinite sum over histories (see section 6.2).
@@
We have seen in section 3.4.5 that tiling problems are very natural candidates to transfer undecidability to physical systems. This was the approach that Kanter (1990) [81] took to prove undecidable properties in a 1D spin system (anisotropic Potts Hamiltonian) using a map to the ’domino snake’ problem, which is a connectivity variation of a Wang tiling problem. The author also indicated how his arguments could be extended to higher dimensions, dynamical evolution, other spin systems, and to some spatial and temporal correlation functions in these systems. He also introduced an uncertainty principle for classical systems caused by undecidability [^20] To our knowledge no further work was done in these directions..
@@
Around the same time, Moore (1990, 1991) [107, 108] showed how to embed a universal Turing Machine into a continuous dynamical system of just three degrees of freedom, corresponding to the motion of a single particle in a finite-sized, three-dimensional box. His construction uses plane and parabolic mirrors in the box to modify the chaotic trajectories of the dynamical system in order to obtain the Turing Machine operations needed for universality. Virtually any question about its long-term behavior is undecidable. We refer to [1, 13] for a discussion of this remarkable achievement.
@@
Taking a very different approach and mapping, da Costa and Doria (1991) [62, 25] (see as well [12]) encode Diophantine equations in classical mechanics and prove undecidability of questions about integrability of certain Hamiltonians and about the existence of chaos in dynamical systems. They rely on results about undecidability of the existence of roots in certain classes of real functions, obtained in turn by mapping to Diophantine equations [109, 110]. But as the authors themselves write, “While we have given general existential proofs concerning undecidability and incompleteness in physics, we have not considered up to now the same questions with respect to particular physically interesting systems …dealing with specific situations is assuredly much more difficult than to prove general facts”. One can find an introduction to these results in [111]. Calude et al. (1995) [16] use the same tools to provide a couple of undecidability results about integral solutions of the one-dimensional heat equation and the electrostatic plane problem.
@@
da Costa et al. (1990) [112] obtained undecidability results in the sense of Gödel in general relativity. They constructed an axiomatic framework for it within Zermelo–Fraenkel set theory and showed that some statements are axiomatically independent. See also [113, 12].
Blondel et al. (2001) [114, 115] proved that several properties of particular classes of discrete time dynamical systems (global convergence, global asymptotic stability, mortality, and nilpotence) are undecidable by simulating Turing Machines with them.
@@
Lloyd (1993, 1994) [116, 117] pioneered fruitful interconnections between quantum information theory and undecidability, which blossomed more than 15 years later, as we will review in detail in the next section. In [116, 117] it is argued that certain spectral properties of quantum mechanical operators are undecidable. The argument is based on the following idea: the unitary evolution $U$ associated to the time evolution of a universal quantum computer has invariant subspaces with discrete spectrum, corresponding to computations that halt, and other invariant subspaces with continuous spectrum, corresponding to computations that do not halt. Via the undecidability of the halting problem, Lloyd concludes that it is undecidable to know if the quantum state associated to a given program (in the infinite-dimensional space in which $U$ is defined) has overlap, or not, with an invariant subspace having a discrete spectrum [^21] See [118, 119] for a discussion on the relation between [116, 117] and [82, 120]..